OpenStudy (anonymous):

given that csc(theta)=7 and theta lies in the second quadrant, find Tan(90degrees-theta)

OpenStudy (anonymous):

does it help to know that \(\tan(90-x)=\cos(x)\)?

OpenStudy (anonymous):

that is wrong sorry, \[\tan(90-x)=\cot(x)\] is what i meant to write

OpenStudy (anonymous):

so the job is to find \(\cot(x)\) knowing \(\csc(x)=7\) we draw a triangle

OpenStudy (anonymous):

that ugly picture is a triangle where the hypotenuse is 7, the opposite side is 1, so the cosecant is 7 we need the third side, which by pythagoras is \(\sqrt{7^2-1^2}=\sqrt{48}\)

OpenStudy (anonymous):

so would that equal Square root -48 for the final answer since its in the second quadrant?

OpenStudy (anonymous):

now we need the cotangent, which is "adjacent over opposite" in other words \[\frac{\sqrt{48}}{1}=\sqrt{48}\] but since you are in quadrant 2 it is negative, so your "final answer" is \[-\sqrt{48}\] yes, what you said

OpenStudy (anonymous):

ok thank you so much this was very helpful

OpenStudy (anonymous):

yw